# Giant electrode effect on tunneling magnetoresistance and electroresistance in van der Waals intrinsic multiferroic tunnel junctions

### v1

@@ Abstract
Van der Waals multiferroic tunnel junctions (vdW-MFTJs) with multiple nonvolatile resistive states are highly suitable for new physics and next-generation storage electronics. However, currently reported vdW-MFTJs are based on two types of materials, i.e., vdW ferromagnetic and ferroelectric materials, forming a multiferroic system. This undoubtedly introduces additional interfaces, increasing the complexity of experimental preparation. Herein, we engineer vdW intrinsic MFTJs utilizing bilayer VS. By employing the nonequilibrium Green’s function combined with density functional theory, we systematically investigate the influence of three types of electrodes (including non-vdW pure metal Ag/Au, vdW metallic 1T-MoS/2H-PtTe, and vdW ferromagnetic metallic FeGaTe/FeGeTe) on the electronic transport properties of VS-based intrinsic MFTJs. We demonstrate that these MFTJs manifest a giant electrode-dependent electronic transport characteristic effect. Comprehensively comparing these electrode pairs, the FGT/FGT electrode combination exhibits optimal transport properties, the maximum TMR (TER) can reach 10949% (69%) and the minimum resistance-area product (RA) is 0.45 ·m, as well as the perfect spin filtering and negative differential resistance effects. More intriguingly, TMR (TER) can be further enhanced to 34000% (380%) by applying an external bias voltage (0.1 V), while RA can be reduced to 0.16 ·m under the influence of biaxial stress (-3%). Additionally, considering the impact of surface dangling bonds of pure metal electrodes on the multiferroicity of VS, we introduce a graphene interlayer between them. This strategy effectively preserves the intrinsic properties of VS and significantly amplifies the TMR (TER) of the MFTJ composed of Ag/Au electrode pairs by an order of magnitude. Our proposed concept of designing vdW-MFTJs using intrinsic multiferroic materials points towards new avenues in experimental exploration.
@@ I Introduction
The concept of vdW-MFTJs was formally introduced by Su et al. [23], based on the FeGeTe/InSe( = 3, 4, 5) heterostructure. Subsequently, Hu et al. [24] proposed /InSe-based MFTJs. Following that, Bai et al. [25] reported low RA in CrSe/CuInPS/CrSe vdW-MFTJ. Then we also designed vdW-MFTJs with six non-volatile resistive states based on FeGeTe/bilayer-InSe/FeGeTe heterostructure [26]. Under the impetus of these studies, vdW-MFTJs have undergone explosive development [27, 28, 29, 30, 31, 32, 33, 34]. However, alongside the flourishing development of vdW-MFTJs, two challenges have emerged that need to be overcome. Firstly, the polarization direction of ferroelectric/ferromagnetic relies entirely on an externally applied electric/magnetic field in MFTJs, resulting in increased energy consumption. Secondly, the currently reported vdW-MFTJs are typically formed by combining ferroelectric and ferromagnetic materials to achieve multiferroicity, undoubtedly introducing more material interfaces and thereby increasing the complexity of experimental preparation. Very recently, the emergence of sliding ferroelectric materials represented by the -BN bilayer [35, 36, 37] and the development of antiferromagnetic tunnel junctions [38, 39, 40] have catered to the former challenge. The polarization direction flip of sliding ferroelectric materials only needs to overcome the weak van der Waals force and the antiferromagnetic tunnel junctions do not require a magnetic pinning layer, and smaller electric and magnetic fields are required. Fortunately, bilayer VS combines intrinsic multiferroicity, interlayer antiferromagnetism, and sliding ferroelectricity all in one [41], making it an ideal candidate material for designing van der Waals intrinsic multiferroic tunnel junctions.
@@
In this work, we theoretically design bilayer VS-based van der Waals intrinsic multiferroic tunnel junctions and investigate the electrode-dependent electronic transport properties by using first-principles computational methods. Here, we select three types of asymmetric electrode pairs, i.e., Ag/Au, 1T-MoS/2H-PtTe, and FeGaTe/FeGeTe. Our calculation results reveal the emergence of non-volatile multiple states with giant TMR and TER, fostered by different types of electrodes, with the optimal electrode option being the van der Waals magnetic FeGaTe/FeGeTe electrode pair. Excitingly, we also observe perfect spin filtering, negative differential resistance effects, and an RA significantly less than 1 ·m (The recording density of 200 Gbit/in requires RA to be less than 1 ·m) within these intrinsic MFTJs. Furthermore, we demonstrate bias- and stress-tunable electron transport properties, with TMR (TER) maximally increased to 34000% (380%) for FeGaTe-VS-FeGeTe MFTJ and RA reduced to 0.015 ·m for Ag-VS-Au MFTJ. Additionally, the insertion of a monolayer graphene not only mitigates the impact of surface dangling bonds on Ag/Au electrodes but also further enhances TMR and TER by an order of magnitude. The design concept of our intrinsic multiferroic tunnel junction holds promise for advancing the development of atomic-scale spintronics devices.
@@ II Computational Methods
The structural relaxation, total energy, and electronic band structure calculations were conducted using density functional theory within the Vienna Ab initio Simulation Package (VASP) [42]. The project-augmented wave pseudopotentials method [43] and the general gradient approximation (GGA) in the Perdew-Burke-Ernzerhof (PBE) were adopted [44]. A cutoff energy of 500 eV and van der Waals correction using the DFT-D3 method were employed [45]. A Monkhorst–Pack -grid [46] was utilized to discretize the Brillouin zone of all VS-based heterojunction systems. In geometry optimization, the convergence criteria for electron energy and force are eV and 0.01 eV/Å, respectively. The PBE+ (on-site Coulomb interaction =3 eV for V atom) was used to treat localized orbitals. Note that the value for other material systems with d orbitals, used as electrodes in this study, was not considered. Ferroelectric polarization is assessed employing the Berry phase method [47]. The spin-polarized quantum transport properties are calculated within the framework of non-equilibrium Green’s functions [48] combined with density functional theory using the Nanodcal software [49]. Also, the GGA with PBE function is employed for the electronic exchange-correlation function in the electron transport calculations. In electronic self-consistent calculations, the cutoff energy was set to 80 Hartree, with a convergence criterion for the Hamiltonian matrix of eV, and the Fermi function temperature was set to 300 K. A -point grid of was employed to calculate current and electronic transmission coefficients.
@@ A. The design of vdW intrinsic MFTJs
Previous study [41] has indicated that the 3-type (ground-state stacking of bilayer VS) stacking bilayer VS exhibits intrinsic multiferroicity, characterized by interlayer antiferromagnetism and spontaneous out-of-plane ferroelectric polarization. Moreover, polarization direction reversal can be achieved through interlayer sliding. Therefore, bilayer multiferroic VS emerges as a candidate material for intrinsic MFTJs. The crystal structures of bilayer VS with two opposite ferroelectric polarization directions are depicted in Fig. 1(a)-(b) and (d)-(e). To further confirm the multiferroicity of bilayer VS, we calculate that the total energy difference between its ferromagnetism and antiferromagnetism is 1.372 meV and its ferroelectric polarization value is determined to be C/m using the Berry phase method [47], which is consistent with the previous study [41]. Fig. 1(b), (e) also include the calculated differential charge density, revealing a distinct difference in charge distribution between the accumulation and depletion regions of the upper and lower layers of VS. This inequivalence in charge results in a net charge transfer between the two layers, and the direction of the transfer reverses with changes in the stacking configuration, giving rise to opposing vertical polarization, which is further corroborated by the marked feature in the plane-averaged differential charge density along the -direction as depicted in the Fig 1(c), (f).
@@
Upon confirming the multiferroicity in bilayer VS, we can construct intrinsic multiferroic tunnel junctions (MFTJs) based on VS. Firstly, we calculate the electronic band structure of the bilayer VS with opposite polarization directions and the corresponding crystal structure is shown in Fig. 2(a), (d). It is observed that, although the ground state of bilayer VS is antiferromagnetic, the band structure exhibits a slight band splitting due to the presence of ferroelectric polarization. Additionally, it is found to be an indirect bandgap semiconductor. Therefore, bilayer VS can serve both as a ferroelectric barrier layer and a magnetic layer, making it suitable for the central scattering region of MFTJs devices. Subsequently, the choice of electrode materials needs to be determined. To meet the requirements of an asymmetric structure on both sides of the central scattering region in the ferroelectric tunnel junction, we select three types of asymmetric electrode pairs: non-vdW pure metal electrodes (Ag/Au), vdW non-magnetic electrodes (1T-MoS/2H-PtTe), and vdW magnetic electrodes (FeGaTe/FeGeTe). In addition to the pure metal electrode pair Ag/Au, Fig. 2(b) and (e), and (c) and (f) illustrate the bulk crystal structures and corresponding electronic band structures of these electrodes, respectively. One can observe that the Fermi level intersects with the entire band structure, indicating metallic properties for these materials and making them suitable for use as electrodes.
@@
After determining the electrodes and the multiferroic barrier layer, we can now construct three types of intrinsic MFTJ devices, namely, Ag/VS/Au, 1T-MoS/VS/2H-PtTe, and FeGaTe/VS/FeGeTe MFTJs. The pure metal Ag/Au(111) has a hexagonal lattice with an in-plane lattice constant of 2.889/2.884 Å in the unit cell. Consistent with the previous studies, the optimized in-plane lattice constants of monolayer 1T-MoS, 2H-PtTe, FeGaTe, FeGeTe, and VS are 3.143 Å[55], 3.895 Å [56], 4.026 Å [57], 4.020 Å [58], and 3.157 Å [59] respectively. Therefore, Ag/Au(111), 1T-MoS, 2H-PtTe, and FeGaTe/FeGeTe have in-plane minimum supercell matches with VS of /(Ag/Au)@(VS), @, @, and /@, respectively. Considering the primary focus on the intrinsic multiferroicity of bilayer VS, the lattice constant of VS is taken as the in-plane lattice constant for all MFTJs as a whole. In this case, the maximum in-plane mismatch ratios obtained for the three types of MFTJs are 5.2%, 6.4%, and 3.75%, respectively. These mismatch ratios are unlikely to occur in realistic experiments due to the weak vdW interactions between layers in the van der Waals systems [13, 31]. However, they must be considered in theoretical simulations due to the periodic boundary conditions. In the following section, we also investigate the influence of in-plane biaxial strain on the transport properties of these MFTJs. Next, we need to determine the optimal stacking configurations at various interfaces of these MFTJs. There are six types of heterojunction interfaces in these MFTJs, i.e., Ag/Au-VS, MoS/PtTe-VS, FeGaTe/FeGeTe-VS. In Ag/Au-VS, four high-symmetry stacking configurations (Ag/Au-bone/hollow/S/V) exist. Fig. 3(c) illustrates the evolution of stacking configurations with the interlayer distance (), revealing that the lowest-energy stacking configuration and interlayer distance are designated as Ag/Au-bone and 2.41/2.535 Å, respectively, with the corresponding optimal structures presented in the inset. For the MoS/PtTe-VS heterojunction, we consider five stacking configurations (V-S/Te, V-hollow, and V-Mo/Pt), as illustrated in Fig. 3(a). The energetically favored stacking configuration is identified as V-S/hollow. Similarly, as shown in Fig. 3(b), one can determine that the optimal stacking arrangement for FeGaTe/FeGeTe-VS is V-Fe/Fe. Finally, based on the stacking sequences of the interfaces established above, we can construct three different intrinsic MFTJs, labeled as Ag-VS-Au, MoS-VS-PtTe, and FeGaTe-VS-FeGeTe, as depicted in Fig. 3(d)-(f). Note that a complete atomic relaxation was performed on the central scattering region with a vacuum layer thickness of 30 Å of these MFTJ devices.
@@ B. Significant TMR/TER and biaxial strain effects at equilibrium
By the physical mechanism of the magnetic/ferroelectric tunnel junction, flipping the magnetization direction of one layer of VS can lead to two opposing magnetic configurations (PC/APC), while sliding the ferroelectric bilayer VS results in two polarization directions. This implies the potential induction of a quadruple resistive state in the VS-based MFTJs. In this section, we first investigate the TMR and TER effects of these three types of intrinsic MFTJs at equilibrium. As presented in Table 1, the transport properties of MFTJs with different electrodes show considerable differences. Distinguished from the other two, the Ag-VS-Au MFTJ displays negative TMR, being -3%/-6% in the ferroelectric PP state. For the MoS-VS-PtTe MFTJ, its TMR is larger than that of Ag-VS-Au MFTJs but much smaller than that of FeGaTe-VS-FeGeTe. Among the three, the highest TMR (7003%/10949% in P/P state) is achieved in the FeGaTe-VS-FeGeTe MFTJ and a perfect spin polarization (SIE=1) in the PC state. In addition to the TMR effect, the TER effect is crucial for assessing the transport performance of MFTJ devices. From Table 1, it is evident that the maximum TER of 153% occurs in the MoS-VS-PtTe MFTJ at APC state, but it is not significantly larger than the other two. We also calculate the resistance-area (RA) product of these MFTJs. It is worth mentioning that the calculated RA products of all four resistance states of Ag-VS-Au MFTJ are less than 0.03 ·m, which is an ideal characteristic of MFTJ for device applications. Additionally, the RA in the FeGaTe-VS-FeGeTe MFTJ is also considerable, being less than 1 ·m. This contrasts sharply with the previously calculated RA products for perovskite-oxide MFTJs of around several ·m [4, 5]. Therefore, overall, the MFTJ composed of vdW magnetic electrodes excels over the other two counterparts due to its giant TMR, larger TER, perfect spin filtering, and RA product less than 1 ·m.
@@
To provide a more detailed insight into the multiple nonvolatile resistance states in these MFTJs, we analyze the Partial Density of States (PDOS) of the central scattering region in the (, ) plane. Here, represents the Fermi energy, and signifies the vertical distance along the transport direction. Figure 4 illustrates the spin-resolved PDOS of the central scattering region in three types of MFTJs, considering opposite ferroelectric polarization and magnetic alignment directions. For the Ag-VS-Au MFTJ, the presence of dangling bonds on the Ag/Au surface forming chemical bonds with VS leads to a modification in the electronic structure of VS. As shown in Fig. 4(a) and (b), there is a significant overlap of wave functions between the bilayer VS at the Fermi level, indicating that this MFTJ does not operate through electron tunneling mechanism, resulting in very low TMR and TER. In Section. D, the insertion of a monolayer graphene between VS and the metal electrodes Ag/Au not only effectively shields the influence of the dangling bonds but also preserves the multiferroic properties of VS. In the other two MFTJs, the little electron density in the deep blue region between the bilayer VS at the Fermi level indicates that electron transport occurs through a tunneling mechanism [see Fig. 4(c)-(f)]. Clearly, a typical TMR effect is evident in the PDOS diagram. We use MoS-VS-PtTe MFTJ as a representative to reveal the TMR effect. Note that here we consider the spin-down DOS of VS at parallel configuration (PC) state as the minority state, as indicated by white circles in Fig. 4(c) and (d). As shown in Fig. 4(c), for the antiparallel configuration (APC) state, electrons with spin-up (spin-down) flow from the left VS with majority (minority) states at the Fermi level, then flow out from the right VS with minority (majority) states. The corresponding relationship between the opposite electron state densities in the bilayer VS implies the high-resistance state. For the PC state, Spin-up electrons are in the majority state in both the upper and lower layers of VS, resulting in low-resistance transport. On the contrary, the VS in the upper and lower layers of spin-down electrons are in minority states near the Fermi level, which hinders electron transport. The above analysis can be directly applied to the FeGaTe-VS-FeGeTe MFTJ.
@@
To further elucidate how ferroelectric polarization and magnetization alignment influence electron transmission, we calculate the -resolved transmission coefficients of these MFTJs at the Fermi level within the 2D Brillouin zone (2D-BZ), as depicted in Fig. 5, which are perpendicular to the transport direction (-axis). As shown in Fig. 5(a)-(d), overall, for the Ag-VS-Au MFTJ, the difference in the ”hot spots” between PC/P and APC/P is small, corresponding to negligible TMR and TER (both less than 10%). The maximum TER (153%) is achieved in the MoS-VS-PtTe MFTJ at the APC state, as clearly revealed in the 2D-BZ electron transmission spectra at the Fermi level. Specifically, by comparing Fig. 5(f) and (h), one can observe that in the P state, the ”hot spots” for spin-up are more concentrated near the high-symmetry points and than in the P state, while the ”hot spots” for spin-down are slightly larger in P than in P. This leads to a significant TER, as indicated by Equation (6). For the FeGaTe-VS-FeGeTe MFTJ, a significantly giant TMR and perfect spin filtering effect can be reflected from the electron transmission coefficients at the Fermi level, with the maximum TMR reaching 10949% in the ferroelectric polarization P state. As depicted in Fig. 5(k) and (l), a substantial number of ”hot spots” emerge in the spin-up channel of the PC state, with almost no ”hot spots” in the spin-down and APC configurations. This suggests the potential presence of significant TMR and high spin polarization rates. Therefore, the distribution of transmission coefficients in the 2D-BZ offers further evidence supporting the pres ence of a giant TMR/TER ratio and a perfect spin-filtering effect in these MFTJs in the equilibrium state.

### v2

@@ Abstract
Van der Waals multiferroic tunnel junctions (vdW-MFTJs) with multiple nonvolatile resistive states are highly suitable for new physics and next-generation storage electronics. However, currently reported vdW-MFTJs are based on two types of materials, i.e., vdW ferromagnetic and ferroelectric materials, forming a multiferroic system. This undoubtedly introduces additional interfaces, increasing the complexity of experimental preparation. Herein, we engineer vdW intrinsic MFTJs utilizing bilayer VS2. By employing the nonequilibrium Green’s function combined with density functional theory, we systematically investigate the influence of three types of electrodes (including non-vdW pure metal Ag/Au, vdW metallic 1T-MoS2/2H-PtTe2, and vdW ferromagnetic metallic Fe3GaTe2/Fe3GeTe2) on the electronic transport properties of VS2-based intrinsic MFTJs. We demonstrate that these MFTJs manifest a giant electrode-dependent electronic transport characteristic effect. Comprehensively comparing these electrode pairs, the Fe3GaTe2/Fe3GeTe2 electrode combination exhibits optimal transport properties, the maximum TMR (TER) can reach 10949% (69%) and the minimum resistance-area product (RA) is 0.45 ·m2, as well as the perfect spin filtering and negative differential resistance effects. More intriguingly, TMR (TER) can be further enhanced to 34000% (380%) by applying an external bias voltage, while RA can be reduced to 0.16 ·m2 under the influence of biaxial stress. Additionally, considering the impact of surface dangling bonds of pure metal electrodes on the multiferroicity of VS2, we introduce a graphene interlayer between them. This strategy effectively preserves the intrinsic properties of VS2 and significantly amplifies the TMR (TER) of the MFTJ composed of Ag/Au electrode pairs by an order of magnitude. Our proposed concept of designing vdW-MFTJs using intrinsic multiferroic materials points towards new avenues in experimental exploration.
@@ I Introduction
The concept of vdW-MFTJs was formally introduced by Su et al. [23], based on the FenGeTe2/In2Se3( = 3, 4, 5) heterostructure. Subsequently, Hu et al. [24] proposed MX2(M=Mn, V, Cr; X=Se, Te)/In2Se3-based MFTJs. Following that, Bai et al. [25] reported low RA in CrSe2/CuInP2S6/CrSe2 vdW-MFTJ. Then we also designed vdW-MFTJs with six non-volatile resistive states based on Fe3GeTe2/bilayer-In2Se3/Fe3GeTe2 heterostructure [26]. Under the impetus of these studies, vdW-MFTJs have undergone explosive development [27, 28, 29, 30, 31, 32, 33, 34]. However, alongside the flourishing development of vdW-MFTJs, two challenges have emerged that need to be overcome. Firstly, the polarization direction of ferroelectric/ferromagnetic relies entirely on an externally applied electric/magnetic field in MFTJs, resulting in increased energy consumption. Secondly, the currently reported vdW-MFTJs are typically formed by combining ferroelectric and ferromagnetic materials to achieve multiferroicity, undoubtedly introducing more material interfaces and thereby increasing the complexity of experimental preparation. Very recently, the emergence of sliding ferroelectric materials represented by the -BN bilayer [35, 36, 37] and the development of antiferromagnetic tunnel junctions [38, 39, 40] have catered to the former challenge. The polarization direction flip of sliding ferroelectric materials only needs to overcome the weak van der Waals force and the antiferromagnetic tunnel junctions do not require a magnetic pinning layer, and smaller electric and magnetic fields are required. Fortunately, bilayer VS2 combines intrinsic multiferroicity, interlayer antiferromagnetism, and sliding ferroelectricity all in one [41], making it an ideal candidate material for designing van der Waals intrinsic multiferroic tunnel junctions.
@@
In this work, we theoretically design bilayer VS2-based van der Waals intrinsic multiferroic tunnel junctions and investigate the electrode-dependent electronic transport properties by using first-principles computational methods. Here, we select three types of asymmetric electrode pairs, i.e., Ag/Au, 1T-MoS2/2H-PtTe2, and Fe3GaTe2/Fe3GeTe2. Our calculation results reveal the emergence of non-volatile multiple states with giant TMR and TER, fostered by different types of electrodes, with the optimal electrode option being the van der Waals magnetic Fe3GaTe2/Fe3GeTe2 electrode pair. Excitingly, we also observe perfect spin filtering, negative differential resistance effects, and an RA significantly less than 1 ·m2 (The recording density of 200 Gbit/in2 requires RA to be less than 1 ·m2) within these intrinsic MFTJs. Furthermore, we demonstrate bias- and stress-tunable electron transport properties, with TMR (TER) maximally increased to 34000% (380%) for Fe3GaTe2-VS2-Fe3GeTe2 MFTJ and RA reduced to 0.015 ·m2 for Ag-VS2-Au MFTJ. Additionally, the insertion of a monolayer graphene not only mitigates the impact of surface dangling bonds on Ag/Au electrodes but also further enhances TMR and TER by an order of magnitude. The design concept of our intrinsic multiferroic tunnel junction holds promise for advancing the development of atomic-scale spintronics devices.
@@ II Computational Methods
The structural relaxation, total energy, and electronic band structure calculations were conducted using density functional theory within the Vienna Ab initio Simulation Package (VASP) [42]. The project-augmented wave pseudopotentials method [43] and the general gradient approximation (GGA) in the Perdew-Burke-Ernzerhof (PBE) were adopted [44]. A cutoff energy of 500 eV and van der Waals correction using the DFT-D3 method were employed [45]. A Monkhorst–Pack -grid [46] was utilized to discretize the Brillouin zone of all VS2-based heterojunction systems. In geometry optimization, the convergence criteria for electron energy and force are eV and 0.01 eV/Å, respectively. The PBE+ (on-site Coulomb interaction eff=3 eV for V atom) was used to treat localized orbitals. Note that the value for other material systems with orbitals, used as electrodes in this study, was not considered. Ferroelectric polarization is assessed employing the Berry phase method [47]. The spin-polarized quantum transport properties are calculated within the framework of non-equilibrium Green’s functions [48] combined with density functional theory using the Nanodcal software [49]. Also, the GGA with PBE function is employed for the electronic exchange-correlation function in the electron transport calculations. In electronic self-consistent calculations, the cutoff energy was set to 80 Hartree, with a convergence criterion for the Hamiltonian matrix of eV, and the Fermi function temperature was set to 300 K. A -point grid of was employed to calculate current and electronic transmission coefficients. Biaxial strain is achieved by directly altering the in-plane lattice constants of the system.
@@ A. The design of vdW intrinsic MFTJs
The previous study [41] has indicated that the 3-type (ground-state stacking of bilayer VS2) stacking bilayer VS2 exhibits intrinsic multiferroicity, characterized by interlayer antiferromagnetism and spontaneous out-of-plane ferroelectric polarization. Moreover, polarization direction reversal can be achieved through interlayer sliding. Therefore, bilayer multiferroic VS2 emerges as a candidate material for intrinsic MFTJs. The crystal structures of bilayer VS2 with two opposite ferroelectric polarization directions are depicted in Fig. 1(a)-(b) and (d)-(e). To further confirm the multiferroicity of bilayer VS2, we calculate that the total energy difference between its ferromagnetism and antiferromagnetism is 1.372 meV and its ferroelectric polarization value is determined to be C/m2 using the Berry phase method [47], which is consistent with the previous study [41]. Fig. 1(b), (e) also include the calculated differential charge density, revealing a distinct difference in charge distribution between the accumulation and depletion regions of the upper and lower layers of VS2. This inequivalence in charge results in a net charge transfer between the two layers, and the direction of the transfer reverses with changes in the stacking configuration, giving rise to opposing vertical polarization, which is further corroborated by the marked feature in the plane-averaged differential charge density along the -direction as depicted in the Fig 1(c), (f).
@@
Upon confirming the multiferroicity in bilayer VS2, we can construct intrinsic multiferroic tunnel junctions (MFTJs) based on VS2. Firstly, we calculate the electronic band structure of the bilayer VS2 with opposite polarization directions and the corresponding crystal structure is shown in Fig. 2(a), (d). It is observed that, although the ground state of bilayer VS2 is antiferromagnetic, the band structure exhibits a slight band splitting due to the presence of ferroelectric polarization [41]. Additionally, it is found to be an indirect bandgap semiconductor. Therefore, bilayer VS2 can serve both as a ferroelectric barrier layer and a magnetic layer, making it suitable for the central scattering region of MFTJs devices. Subsequently, the choice of electrode materials needs to be determined. To meet the requirements of an asymmetric structure on both sides of the central scattering region in the ferroelectric tunnel junction, we select three types of asymmetric electrode pairs: non-vdW pure metal electrodes (Ag/Au), vdW non-magnetic electrodes (1T-MoS2/2H-PtTe2), and vdW magnetic electrodes (Fe3GaTe2/Fe3GeTe2). In addition to the pure metal electrode pair Ag/Au, Fig. 2(b) and (e), and (c) and (f) illustrate the bulk crystal structures and corresponding electronic band structures of these electrodes, respectively. One can observe that the Fermi level intersects with the entire band structure, indicating metallic properties for these materials and making them suitable for use as electrodes.
@@
After determining the electrodes and the multiferroic barrier layer, we can now construct three types of intrinsic MFTJ devices, namely, Ag/VS2/Au, 1T-MoS2/VS2/2H-PtTe2, and Fe3GaTe2/VS2/Fe3GeTe2 MFTJs. The pure metal Ag/Au(111) has a hexagonal lattice with an in-plane lattice constant of 2.889/2.884 Å in the unit cell. Consistent with the previous studies, the optimized in-plane lattice constants of monolayer 1T-MoS2, 2H-PtTe2, Fe3GaTe2, Fe3GeTe2, and VS2 are 3.143 Å[55], 3.895 Å [56], 4.026 Å [57], 4.020 Å [58], and 3.157 Å [59] respectively. Therefore, Ag/Au(111), 1T-MoS2, 2H-PtTe2, and Fe3GaTe2/Fe3GeTe2 have in-plane minimum supercell matches with VS2 of /(Ag/Au)@(VS2), @, @, and /@, respectively. Considering the primary focus on the intrinsic multiferroicity of bilayer VS2, the lattice constant of VS2 is taken as the in-plane lattice constant for all MFTJs as a whole. In this case, the maximum in-plane mismatch ratios obtained for the three types of MFTJs are 5.2%, 6.4%, and 3.75%, respectively. These mismatch ratios are unlikely to occur in realistic experiments due to the weak vdW interactions between layers in the van der Waals systems [13, 31]. However, they must be considered in theoretical simulations due to the periodic boundary conditions. In the following section, we also investigate the influence of in-plane biaxial strain on the transport properties of these MFTJs. Next, we need to determine the optimal stacking configurations at various interfaces of these MFTJs. There are six types of heterojunction interfaces in these MFTJs, i.e., Ag/Au-VS2, MoS2/PtTe2-VS2, Fe3GaTe2/Fe3GeTe2-VS2. In Ag/Au-VS2, four high-symmetry stacking configurations (Ag/Au-bone/hollow/S/V) exist. Fig. 3(c) illustrates the evolution of stacking configurations with the interlayer distance (), revealing that the lowest-energy stacking configuration and interlayer distance are designated as Ag/Au-bone and 2.41/2.535 Å, respectively, with the corresponding optimal structures presented in the inset. For the MoS2/PtTe2-VS2 heterojunction, we consider five stacking configurations (V-S/Te, V-hollow, and V-Mo/Pt), as illustrated in Fig. 3(a). The energetically favored stacking configuration is identified as V-S/hollow. Similarly, as shown in Fig. 3(b), one can determine that the optimal stacking arrangement for Fe3GaTe2/Fe3GeTe2-VS2 is V-Fe/Fe. Finally, based on the stacking sequences of the interfaces established above, we can construct three different intrinsic MFTJs, labeled as Ag-VS2-Au, MoS2-VS2-PtTe2, and Fe3GaTe2-VS2-Fe3GeTe2, as depicted in Fig. 3(d)-(f). Note that a complete atomic relaxation was performed on the central scattering region with a vacuum layer thickness of 30 Å of these MFTJ devices.
@@ B. Significant TMR/TER and biaxial strain effects at equilibrium
By the physical mechanism of the magnetic/ferroelectric tunnel junction, flipping the magnetization direction of one layer of VS2 can lead to two opposing magnetic configurations (PC/APC), while sliding the ferroelectric bilayer VS2 results in two polarization directions. This implies the potential induction of a quadruple resistive state in the VS2-based MFTJs. In this section, we first investigate the TMR and TER effects of these three types of intrinsic MFTJs at equilibrium. As presented in Table 1, the transport properties of MFTJs with different electrodes show considerable differences. Distinguished from the other two, the Ag-VS2-Au MFTJ displays negative TMR, being -3%/-6% in the ferroelectric PP state. For the MoS2-VS2-PtTe2 MFTJ, its TMR is larger than that of Ag-VS2-Au MFTJs but much smaller than that of Fe3GaTe2-VS2-Fe3GeTe2. Among the three, the highest TMR (7003%/10949% in P/P state) is achieved in the Fe3GaTe2-VS2-Fe3GeTe2 MFTJ and a perfect spin polarization (SIE=1) in the PC state. In addition to the TMR effect, the TER effect is crucial for assessing the transport performance of MFTJ devices. From Table 1, it is evident that the maximum TER of 153% occurs in the MoS2-VS2-PtTe2 MFTJ at APC state, but it is not significantly larger than the other two. We also calculate the resistance-area (RA) product of these MFTJs. It is worth mentioning that the calculated RA products of all four resistance states of Ag-VS2-Au MFTJ are less than 0.03 ·m2, which is an ideal characteristic of MFTJ for device applications. Additionally, the RA in the Fe3GaTe2-VS2-Fe3GeTe2 MFTJ is also considerable, being less than 1 ·m2. This contrasts sharply with the previously calculated RA products for perovskite-oxide MFTJs of around several ·m2 [4, 5]. Therefore, overall, the MFTJ composed of vdW magnetic electrodes excels over the other two counterparts due to its giant TMR, larger TER, perfect spin filtering, and RA product less than 1 ·m2.
@@
To provide a more detailed insight into the multiple nonvolatile resistance states in these MFTJs, we analyze the Partial Density of States (PDOS) of the central scattering region in the (, ) plane. Here, represents the Fermi energy, and signifies the vertical distance along the transport direction. Figure 4 illustrates the spin-resolved PDOS of the central scattering region in three types of MFTJs, considering opposite ferroelectric polarization and magnetic alignment directions. For the Ag-VS2-Au MFTJ, the presence of dangling bonds on the Ag/Au surface forming chemical bonds with VS2 leads to a modification in the electronic structure of VS2. As shown in Fig. 4(a) and (b), there is a significant overlap of wave functions between the bilayer VS2 at the Fermi level, indicating that this MFTJ does not operate through electron tunneling mechanism, resulting in very low TMR and TER. In Section. D, the insertion of a monolayer graphene between VS2 and the metal electrodes Ag/Au not only effectively shields the influence of the dangling bonds but also preserves the multiferroic properties of VS2. In the other two MFTJs, the little electron density in the deep blue region between the bilayer VS2 at the Fermi level indicates that electron transport occurs through a tunneling mechanism [see Fig. 4(c)-(f)]. Clearly, a typical TMR effect is evident in the PDOS diagram. We use MoS2-VS2-PtTe2 MFTJ as a representative to reveal the TMR effect. Note that here we consider the spin-down DOS of VS2 at parallel configuration (PC) state as the minority state, as indicated by white circles in Fig. 4(c) and (d). As shown in Fig. 4(c), for the antiparallel configuration (APC) state, electrons with spin-up (spin-down) flow from the left VS2 with majority (minority) states at the Fermi level, then flow out from the right VS2 with minority (majority) states. The corresponding relationship between the opposite electron state densities in the bilayer VS2 implies the high-resistance state. For the PC state, spin-up electrons are in the majority state in both the upper and lower layers of VS2, resulting in low-resistance transport. On the contrary, the VS2 in the upper and lower layers of spin-down electrons are in minority states near the Fermi level, which hinders electron transport. The above analysis can be directly applied to the Fe3GaTe2-VS2-Fe3GeTe2 MFTJ.
@@
To further elucidate how ferroelectric polarization and magnetization alignment influence electron transmission, we calculate the -resolved transmission coefficients of these MFTJs at the Fermi level within the 2D Brillouin zone (2D-BZ), as depicted in Fig. 5, which are perpendicular to the transport direction (-axis). As shown in Fig. 5(a)-(d), overall, for the Ag-VS2-Au MFTJ, the difference in the ”hot spots” between PC/P and APC/P is small, corresponding to negligible TMR and TER (both less than 10%). The maximum TER (153%) is achieved in the MoS2-VS2-PtTe2 MFTJ at the APC state, as clearly revealed in the 2D-BZ electron transmission spectra at the Fermi level. Specifically, by comparing Fig. 5(f) and (h), one can observe that in the P state, the ”hot spots” for spin-up are more concentrated near the high-symmetry points and than in the P state, while the ”hot spots” for spin-down are slightly larger in P than in P. This leads to a significant TER, as indicated by Equation (6). For the Fe3GaTe2-VS2-Fe3GeTe2 MFTJ, a significantly giant TMR and perfect spin filtering effect can be reflected from the electron transmission coefficients at the Fermi level, with the maximum TMR reaching 10949% in the ferroelectric polarization P state. As depicted in Fig. 5(k) and (l), a substantial number of ”hot spots” emerge in the spin-up channel of the PC state, with almost no ”hot spots” in the spin-down and APC configurations. This suggests the potential presence of significant TMR and high spin polarization rates. Therefore, the distribution of transmission coefficients in the 2D-BZ offers further evidence supporting the pres ence of a giant TMR/TER ratio and a perfect spin-filtering effect in these MFTJs in the equilibrium state.